# Negative feedback and oscillations in a model for mRNA translation

### v1

@@ Acknowledgments
We thank Alexander Ovseevich for helpful comments.
V Appendix: Proofs
@@
Proof of Proposition 2. Since $[0,1]^{n}$ is convex and compact, it follows from Proposition 1 and Brouwer’s fixed point theorem [22] that the closed-loop system admits at least one equilibrium $e\in[0,1]^{n}$. Using Proposition 1 implies that in fact $e\in(0,1)^{n}$. By (II), the equilibrium satisfies
@@
These equations imply in particular that $e_{n}$ uniquely determines $e_{i}$ for all $i$. In order to prove uniqueness of the equilibrium, seeking a contradiction, suppose that $p,q\in(0,1)^{n}$ are two different equilibrium points. Then $p_{n}\not=q_{n}$, and we may assume that $p_{n}>q_{n}$. By (V), $e_{n-1}=\frac{\lambda_{n}e_{n}}{\lambda_{n-1}(1-e_{n})}$, so $p_{n-1}>q_{n-1}$. Eq. (V) also gives $e_{n-2}=\frac{\lambda_{n}e_{n}}{\lambda_{n-2}(1-e_{n-1})}$, so $p_{n-2}>q_{n-2}$. Continuing in this way, we find that $p_{i}>q_{i}$ for all $i$. Now (V) and the fact that $g$ is decreasing give
@@
Let $s_{i}(x)$ denote the sum of the entries in column $i$ of $J(x)$, with the off-diagonal terms taken with absolute value. Then
@@
Proof of Theorem 3. Since $g^{\prime}(z)<0$ for all $z\geq 0$, Eq. (7) implies that $J(x)$ has the sign pattern of a 2-cooperative system [46] for all $x\in[0,1]^{n}$, and it is irreducible for all $x\in(0,1)^{n}$. Using Theorem 2 in [21] completes the proof of Theorem 3.

### v2

@@ Acknowledgments
We thank Alexander Ovseevich and Tamir Tuller for helpful comments. We are grateful to the anonymous reviewers and the editor for many helpful comments that helped us to improve this paper.
Appendix 1: Proofs
@@
Proof of Proposition 2. Since $0,1]^{n}$ is convex and compact, it follows from Proposition 1 and Brouwer’s fixed point theorem [30] that the closed-loop system admits at least one equilibrium $e\in[0,1]^{n}$. Using Proposition 1 implies that in fact $e\in(0,1)^{n}$. By (II), the equilibrium satisfies
@@
These equations imply in particular that $e_{n}$ uniquely determines $e_{i}$ for all $i$. In order to prove uniqueness of the equilibrium, seeking a contradiction, suppose that $p,q\in(0,1)^{n}$ are two different equilibrium points. Then $p_{n}\not=q_{n}$, and we may assume that $p_{n}>q_{n}$. By (Appendix 1: Proofs), $e_{n-1}=\frac{\lambda_{n}e_{n}}{\lambda_{n-1}(1-e_{n})}$, so $p_{n-1}>q_{n-1}$. Eq. (Appendix 1: Proofs) also gives $e_{n-2}=\frac{\lambda_{n}e_{n}}{\lambda_{n-2}(1-e_{n-1})}$, so $p_{n-2}>q_{n-2}$. Continuing in this way, we find that $p_{i}>q_{i}$ for all $i$. Now (Appendix 1: Proofs) and the fact that $g$ is decreasing give
@@
Recall that $g^{\prime}(z)<0$ for any $z>0$, so entry $(1,n)$ in $J(x)$ is non-positive. Let $s_{i}(x)$ denote the sum of the entries in column $i$ of $J(x)$, with the off-diagonal terms taken with absolute value. Then
@@
Proof of Theorem 3. A 2-cooperative dynamic system has a Jacobian with the following sign pattern
$\bar{A}_{2}:=\left[\begin{smallmatrix}*&\geq 0&0&\dots&0&\leq 0\\ \geq 0&*&\geq 0&\dots&0&0\\ 0&\geq 0&*&\dots&0&0\\ &&\vdots\\ 0&0&0&\dots&*&\geq 0\\ \leq 0&0&0&\dots&\geq 0&*\\ \end{smallmatrix}\right]$ (10)
where ”*” means that the sign of the entry is not important. Since $g^{\prime}(z)<0$ for all $z\geq 0$, Eq. (9) implies that $J(x)$ has the sign pattern of a 2-cooperative system [56] for all $x\in[0,1]^{n}$, and it is irreducible for all $x\in(0,1)^{n}$. Using Theorem 2 in [29] completes the proof of Theorem 3.
Appendix 2: Matlab code for generating Example 1
%%%%% file negtive_RFM
global k  % controller gain
global lam0 lam1 lam2 lam3 % transition rates along the chain
lam0=12   ;lam1 =6 ; lam2=5 ; lam3= 8  ;
set(gcf,’renderer’,’painters’)
figure (1); hold off;
k=0.1; % low gain
init=[0 0 0]’;plot3(0,0,0,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
init=[1/2 1/2 1/2]’; plot3(1/2,1/2,1/2,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
init=[.1 .4 .4]’;plot3(.1,.4,.4,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
init=[.4 .2 .2]’;plot3(.4,.2,.2,’ok’);hold on;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);
box on;
grid on;
plot3(0.1855, 0.2004, 0.1113,’*k’); % plot the eq
 xlabel(’\$x_1\$’,Interpreter=’latex’,FontSize=18);
 ylabel(’\$x_2\$’,Interpreter=’latex’,FontSize=18);
 zlabel(’\$x_3\$’,Interpreter=’latex’,FontSize=18);
set(gcf,’renderer’,’painters’)
figure(2); hold off;
k=1;   %high gain
init=[0 0 0]’;
[t,x]=ode45(@dyn_negative_RFM,[0 100],init);
plot3(x(:,1),x(:,2),x(:,3),’k’);hold on;
plot3(0.2472 ,0.2560 ,0.1380 ,’*k’); % plot the eq
box on
grid on
xlabel(’\$x_1\$’,Interpreter=’latex’,FontSize=18);
ylabel(’\$x_2\$’,Interpreter=’latex’,FontSize=18);
zlabel(’\$x_3\$’,Interpreter=’latex’,FontSize=18);
function ret=dyn_negative_RFM(t,x)
global k   % controller gain
global lam0 lam1 lam2 lam3 % transition rates along the chain
x1=x(1);x2=x(2);x3=x(3);
u= k*1/(1+(lam3*x3)^20 ); % the negative feedback
x1d = lam0*u*(1-x1)-lam1*x1*(1-x2);
x2d = lam1*x1*(1-x2)-lam2*x2*(1-x3);
x3d = lam2*x2*(1-x3)-lam3*x3;
ret=[x1d;x2d;x3d];
Appendix 3: Matlab code for calculating the Jacobian in Example 2
lam0=0.1678;
lam1=0.2572;
lam2=0.2758;
lam3=0.2514;
lam4=0.2612;
lam5=0.3002;
k=10;
a=100;
%% The eq equations
fun = @(x) [lam0*k* (1-tanh(a*lam5*x(5)))*(1-x(1))-lam1*x(1)*(1-x(2));
lam1*x(1)*(1-x(2))-lam2*x(2)*(1-x(3));
lam2*x(2)*(1-x(3))-lam3*x(3)*(1-x(4));
lam3*x(3)*(1-x(4))-lam4*x(4)*(1-x(5));
lam4*x(4)*(1-x(5))-lam5*x(5) ];
x0 = [1/2; 1/2; 1/2; 1/2; 1/2];% an initial guess
[x, fval, exitflag, output] = fsolve(fun, x0,options);
e=x; %%%   eq point
% calculate the Jacobain J(e)
y=lam5*e(5);
g=1-tanh(a*y) ;
gdot= -a/ (cosh(a*y))^2 ;
J=[-lam0*k*g-lam1*(1-e(2))  lam1*e(1)  0  0  lam0*lam5*(1-e(1))*k*gdot ;
   lam1*(1-e(2))  -lam1*e(1)-lam2*(1-e(3)) lam2*e(2)   0             0;
    0           lam2*(1-e(3))   -lam2*e(2)-lam3*(1-e(4)) lam3*e(3)   0;
    0     0      lam3*(1-e(4))     -lam3*e(3)-lam4*(1-e(5)) lam4*e(4) ;
    0     0             0           lam4*(1-e(5)) -lam4*e(4)-lam5    ];
eig(J)